OpenStudy (anonymous):

Please check 1.p(x) =(x-3)+2 Ans: Translation to the right 3 units. 2.) Translation down 1 unit and to the left 3 units. Ans: g(x) =(x+5)^2 =1 3.) F(x)=3x +5; x axis Ans: F(-x)=3(x)+5 = 3x+5 4.) Vertical stretch by a factor of 2. . . .I don't know?

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

one is correct

OpenStudy (anonymous):

Yay^^

OpenStudy (gorv):

second me the Q

OpenStudy (gorv):

send me 2nd Q

OpenStudy (anonymous):

OOps, Write an equation for each transformation of y=x^2

OpenStudy (gorv):

okkk so now new value y-(-1)=(x+3)^2

OpenStudy (gorv):

down one unit means y=-1

OpenStudy (anonymous):

Ah. . . It's a bit perplexing at times, x3

OpenStudy (gorv):

x shift left ward so x+3

OpenStudy (e.mccormick):

A vertical stretch means all up and down values are larger. This one says by 2. So you would put somrthing in there to multiply the answers by 2.

OpenStudy (anonymous):

g(x) = 2(2x-7) =. . .? xD

OpenStudy (e.mccormick):

Here are two references on transformations that have some nice diagrams to help show what happens. http://www.mathsisfun.com/sets/function-transformations.html http://www.regentsprep.org/Regents/math/algtrig/ATP9/funclesson1.htm

OpenStudy (e.mccormick):

And yes, that looks like a vertical stretch.

OpenStudy (anonymous):

Thanks! Sorry I couldn't medal you.

OpenStudy (anonymous):

@e.mccormick g(x) = 2(2x-7) =4x-14?